OpenStudy (breathlessreaper):

The function below shows the number of students of a school who enrolled for cooking classes. Let f(x) represent the total number of students who enrolled for the classes after x years: f(x) = 11(1.35)x The average rate of change in the number of students who enrolled for cooking classes from the first to the fifth year is ________students per year. Round your answer to the nearest whole number.

OpenStudy (anonymous):

might this be \[11\times (1.35)^x\]?

OpenStudy (breathlessreaper):

for my answer i had gotten 45 when i did it, not sure how that would work... cause i believe it was supposed to be something like solve for f(x) five times (with 1, 2, 3, 4, and 5) and then find the average rate of change

OpenStudy (breathlessreaper):

and to do that you just see how much they changed each year, add them all up, and divide by 5.... @satellite73

OpenStudy (anonymous):

no i think you only need \[\frac{f(5)-f(1)}{5-1}\]

OpenStudy (anonymous):

my only problem is that usually "first year" means that \(x=0\) and "fifth year" makes \(x=4\) so perhaps it should be \[\frac{f(4)-f(0)}{4}\]

OpenStudy (anonymous):

i understand why you thought that, but "average rate of change" does not mean the "average of the yearly rates of change" it just means the average over the whole time period

OpenStudy (breathlessreaper):

I'm just guessing, I don't really understand how to get it o-o

OpenStudy (anonymous):

first off, it is really \[f(x)=11\times (1.35)^x\]?

OpenStudy (anonymous):

i would compute \[\frac{11\times (1.35)^4-11}{4}\] and round it to 6 http://www.wolframalpha.com/input/?i=%2811%281.35%29^4-11%29%2F4

OpenStudy (breathlessreaper):

Yeppp

OpenStudy (anonymous):

yeppp as in "that is the function" or yeppp as in "that is the right answer"?

OpenStudy (breathlessreaper):

Yeppp as in thats the function. @satellite73

OpenStudy (breathlessreaper):

Sorry about my slow internet...